OpenStudy (anonymous):

find all the zeros of: f(x)= x^3+5x^2-8x-40

OpenStudy (anonymous):

given \[40=2^45\] the possible zeroes are: \[\pm 1,\pm 2,\pm 4, \pm 5, \pm 10, \pm 16, \pm 20, \pm 40\] because it is -40 not +40 either 1 or 3 zeroes must be negative. because we have +5x^2 not - 5x^2 at least one must be positive. thise means our answer will look like: \[(x+a)(x+b)(x-c)\] 5=a+b-c -8=ab-ac-bc -40=abc

OpenStudy (anonymous):

where did you get 40=2^4(5)

OpenStudy (anonymous):

i'm sorry that should be that should be 2^3(5)

OpenStudy (anonymous):

oh ok that makes better sense

OpenStudy (anonymous):

so would the answer be -5,+or- 2sqrt2

OpenStudy (anonymous):

not quite. that would leave us with a positive 40

OpenStudy (anonymous):

only one number can be negative.

OpenStudy (anonymous):

that's strange because that's what my teacher put for the answer. I just didn't know how she got it

OpenStudy (anonymous):

i'm sorry. my mistake that is correct.

OpenStudy (anonymous):

\[-5,\pm2\sqrt{2}\]

OpenStudy (anonymous):

yes i had misread the previous one.

OpenStudy (anonymous):

in order to reach that you would have to use synthetic division to find the first possible zero.

OpenStudy (anonymous):

ok thank you for the help :)

OpenStudy (anonymous):

because we have 5x^2 i would suggest 5 as the first possibility

OpenStudy (anonymous):

of course :) so synthetic division would leave \[(x+5)(x^2-8)\] can you solve it from here?

OpenStudy (anonymous):

of course thanks